OpenStudy (anonymous):

find the exact trigonometric function value sec 495º

OpenStudy (anonymous):

Helpp!

OpenStudy (anonymous):

495 degrees = (360 + 135) degrees 360 is whole circle so we can just ignore that and focus on 135. 135 degrees = (90 + 45) degrees 90 degrees = pi/2 and 45 degrees = pi/4 Try to get the rest done yourself ;)